# Order Is All You Need for Categorical Data Clustering

### v1

@@
Abstract.
Categorical data composed of nominal valued attributes are ubiquitous in knowledge discovery and data mining tasks. Due to the lack of well-defined metric space, categorical data distributions are difficult to intuitively understand. Clustering is a popular technique suitable for data analysis. However, the success of clustering often relies on reasonable distance metrics, which happens to be what categorical data naturally lack. Therefore, the cluster analysis of categorical data is considered a critical but challenging problem. This paper introduces the new finding that the order relation among attribute values is the decisive factor in clustering accuracy, and is also the key to understanding the categorical data clusters. To automatically obtain the orders, we propose a new learning paradigm that allows joint learning of clusters and the orders. It turns out that clustering with order learning achieves superior clustering accuracy, and the learned orders provide intuition for understanding the cluster distribution of categorical data. Extensive experiments with statistical evidence and case studies have verified the effectiveness of the new “order is all you need” insight and the proposed method. The source code of this submission is temporarily opened [^1] https://anonymous.4open.science/r/OCL-demo.
@@ Abstract.
submissionid: 51ccs: Computing methodologies Cluster analysisccs: Computing methodologies Learning latent representationsccs: Computing methodologies Mixture modeling
1. Introduction
@@
In the literature, there are two main types of clustering algorithms for categorical data: 1) $k$-modes and its subsequent variants (Huang, 1998)(Bai et al., 2011)(Jia and Cheung, 2018) that partition data samples into clusters based on certain distance measures, and 2) cluster quality assessment-based algorithms (Hu et al., 2022)(Li et al., 2004) that adjust the sample-cluster affiliations based on certain cluster quality measures. Their performance is commonly dominated by the metric space represented by the adopted measures. Thus, the focus of subsequent research has shifted from designing clustering algorithms to proposing reasonable measures that suit the categorical data clustering (Naouali et al., 2020).
@@
The conventional Hamming distance measure and one-hot encoding strategy are very common in categorical data clustering, but they both simply represent the distance between two attribute values based on whether the two values are identical or not, which is far from informative. Later, on the one hand, more advanced distance measures have been proposed to discriminatively indicate the value-level differences by considering the intra-attribute probability of possible values (Sandler and Lindenbaum, 2011), inter-attribute relationship (Ahmad and Dey, 2007)(Le and Ho, 2005)(Ienco et al., 2012), or both (Jia et al., 2016)(Zhang and Cheung, 2022b). On the other hand, advanced representation strategies that encode distances among data samples (Qian et al., 2015), and couplings in the attribute and value levels (Jian et al., 2018c)(Jian et al., 2017), have also been proposed Although all the above advanced techniques have considerably improved the clustering performance, they are all task-independent and thus incompetent in adapting different clustering tasks on various categorical datasets.
[…]
@@
It is noteworthy that order is the actual basis for clustering in most cases. For instance, when a $k$-means-type algorithm tries to partition a sample into one of the $k$ clusters, it actually obtains the overall order of the $k$ clusters in terms of admitting the sample. The overall order is formed based on the similarity orders of the clusters to the sample reflected by different attributes. Note that each numerical attribute corresponds to a one-dimensional Euclidean distance space with a clear order of values, which makes it easy to obtain the value order. According to the above analysis, we also note that although most existing categorical data clustering approaches do not specify the order of attribute values, their clustering process still passively forms the order through similarity measurement, regardless of whether there is an obvious semantic order of the attribute values.
@@
Therefore, knowing the correct order of attribute values may be beneficial to categorical data clustering, and the experimental results in Figure 1 confirmed this. It can be seen that even randomly generated orders can easily achieve significant accuracy improvements. Another interesting observation is that, although attribute values of some datasets (i.e., the right eight datasets) are with obvious semantic orders, they are rarely optimal for clustering. This is because the semantic orders of attributes are relatively independent of specific clustering tasks, and this information gap can prevent some semantic orders from well-serving clustering. Therefore, obtaining the optimal orders w.r.t. certain clustering tasks is the key to achieving accurate categorical data clustering.
@@
This paper, therefore, proposes a new clustering paradigm that dynamically learns orders based on the current data partition during clustering. Since the order learning is connected to the clustering task, the optimal order for clustering can thus be automatically obtained. To make full use of the learned orders, a distance learning mechanism is also introduced to dynamically measure the sample-cluster affiliation under the guidance of the orders. It turns out that the learned orders can significantly improve clustering accuracy and can also intuitively help understand the distribution and clustering effect of categorical data. Extensive experimental evaluations demonstrate the superiority of the proposed method. The main contributions can be summarized into three points:
@@
  • A new insight that “obtaining the optimal orders is the decisive factor for accurately clustering categorical data” is proposed for the first time. Accordingly, a new clustering paradigm with order distance learning is proposed.
@@
  • We propose a distance structure learning mechanism to automatically obtain the order and distance. This learning process is jointly performed with clustering, making the clustering performance more robust to various datasets.
@@
  • Both the learning process and the learned order are highly interpretable, which is helpful for understanding complex categorical data. Moreover, the algorithm is efficient, strictly convergent, and can be easily extended to mixed data [^2] Dataset composed of a mixture of categorical and numerical attributes..
2. Related Work
@@
This section overviews the related works in data encoding, representation learning, and consideration of value order, for categorical data clustering.
@@
The data encoding techniques convert the categorical data into numerical values based on certain strategies, and then perform clustering on the numerical encoding. The conventional one-hot encoding and Hamming distance construct similar distance structures for categorical data by assigning distances “1” and “0” to any pair of different and identical values, respectively. As they cannot distinguish different dissimilarity degrees (dos Santos and Zárate, 2015)(Boriah et al., 2008) of data samples, information loss is usually unavoidable during clustering. Some statistic-based distance measures (Lin, 1998)(Cheung and Jia, 2013)(dos Santos and Zárate, 2015) consider the intra-attribute occurrence frequency of values to better distinguish the data samples. They take a common principle that two similar values may have more consistent statistics within the corresponding attribute or cluster. Some metrics (Ahmad and Dey, 2007)(Le and Ho, 2005) further define distance considering the relevance between attributes. The representative measure proposed in (Ienco et al., 2012) selects more relevant attributes for the distance measurement. Later, the measures (Jia et al., 2016)(Jian et al., 2018a) that more comprehensively consider both the intra- and inter-attribute couplings reflected by the data statistics have been presented. Coupling-based encoding strategies (Jian et al., 2018c)(Jian et al., 2017) have also been designed accordingly to encode the couplings. All the above techniques perform value-level encoding to reflect sample-level distances, while the work (Qian et al., 2015) directly encodes a data sample by concatenating the distances from it to all the other samples.
@@
Recently, more advanced representation learning methods have been proposed to learn the categorical data encoding. The work proposed in (Zhu et al., 2022) projects dataset through different kernels for more informative coupling representation, and then learns the kernel parameters to obtain task-oriented representation. Some approaches (Jian et al., 2018b)(Zhu et al., 2020) have also been designed to simultaneously encode the numerical and categorical attributes. However, the involved hyper-parameters have a great impact on their performance, and the appropriate parameter selection will be laborious in different real applications. Although a parameter-free learning approach (Jia and Cheung, 2018) has been proposed to learn sample-cluster distance and number of clusters simultaneously, the distance is updated dependent on cluster statistics and thus cannot reveal the natural distance structure of categorical data.
@@
Some new advances proposed with the consideration of value order further distinguish categorical attributes into nominal and ordinal attributes. The distance metrics (Zhang et al., 2020)(Zhang and Cheung, 2022b) define the distance between attribute values based on the data statistical information with considering the semantic order. Later, a representation learning approach was introduced by (Zhang and Cheung, 2020), which treats the possible values of an ordinal attribute as linearly arranged concepts, and learns the distances between adjacent concepts. Later, the works (Zhang and Cheung, 2022a)(Zhang and Cheung, 2023) further unify the distances of nominal and ordinal attributes, and make them learnable with clustering. The most recent work (Zhang et al., 2022) attempts to convert the nominal attributes into ordinal ones through geometric projection, and then learn the latent distance spaces during clustering. The above recent advances achieve considerable clustering performance improvements and empirically prove the importance of value order in categorical data clustering. However, they all assume the availability of semantic order of attribute values. Moreover, they directly adopt this order for clustering, and overlook that the semantic order may not be suitable under different clustering scenarios as shown in Figure 1.
3. Proposed Method
@@
This section first formulates the problem, then proposes the order distance definition and order learning mechanism, and finally introduces the joint learning of the orders and clusters. Analyses of time complexity and convergence are also provided.
3.1. Problem Formulation
The problem of categorical data clustering with order learning is formulated as follows. Given a categorical dataset $X=\{\mathbf{x}_{1},\mathbf{x}_{2},...,\mathbf{x}_{n}\}$ with $n$ data samples (e.g., a collection of $n$ bank clients). Each data sample $\mathbf{x}_{i}$ can be denoted as a $d$-dimensional row vector $\mathbf{x}_{i}=[x_{i,1},x_{i,2},..,x_{i,d}]^{\top}$ represented by $d$ attributes $A=\{\mathbf{a}_{1},\mathbf{a}_{2},...,\mathbf{a}_{d}\}$ (e.g., occupation, gender, priority, etc.). Each attribute $\mathbf{a}_{r}$ (e.g., occupation) can be denoted as a column vector $\mathbf{a}_{r}=[x_{1,r},x_{2,r},...,x_{n,r}]$ composed of the occupation description of all the $n$ clients. For the categorical attribute “occupation” $\mathbf{a}_{r}$, all its $n$ values are taken from a limited number of qualitative possible values $V_{r}=\{v_{r,1},v_{r,2},...,v_{r,l_{r}}\}$, e.g., $\{\text{lawyer},\text{doctor},...,\text{scientist}\}$. The subscripts indicate the sequential number of an attribute and the sequential number of the attribute’s possible value. For example, doctor is the second value in the $r$-th attribute “occupation”, so we have $v_{r,2}=\text{doctor}$. We use $l_{r}$ to indicate the total number of possible values of an attribute, and it usually satisfies $l_{r}\ll n$ for real categorical data. Please note that the sequential numbers of possible values only distinguish them, and do not indicate their order.
  1. (1)
    $\text{dist}(\text{doctor},\text{scientist})>\text{dist}(\text{doctor},\text{lawyer})$,
  2. (2)
    $\text{dist}(\text{doctor},\text{scientist})>\text{dist}(\text{lawyer},\text{scientist})$,
because “doctor” and “scientist” are further away in order, compared to “lawyer” and “doctor” that are adjacent in order.
We aim to approximate the optimal orders of all the $d$ attributes $O^{*}=\{O^{*}_{1},O^{*}_{2},...,O^{*}_{d}\}$. Then the order information reflecting the distance structure can guide the partition of data objects during clustering. Our approach, i.e., learning order relations, is more generative than most existing approaches that directly define the distance between attribute values. We focus more on how the intrinsic relationship between values affects the distance between samples in forming clusters, rather than simply defining the clustering-irrelevant distance between values based on data statistics, e.g., occurrence probability of values in an attribute.
@@ 3.1. Problem Formulation
For a crisp partitional clustering task to partition $X$ into $k$ non-overlapping subsets $C=\{C_{1},C_{2},...,C_{k}\}$, the objective can be formalized as the problem of minimizing:
(1) $L(\mathbf{Q},O)=\sum_{m=1}^{k}\sum_{i=1}^{n}q_{i,m}\cdot\Theta(\mathbf{x}_{i},C_{m};O)$
@@
(2) $q_{i,m}=\left\{\begin{array}[]{ll}1,&\text{if}\ m=\arg\min\limits_{y}\Theta(\mathbf{x}_{i},C_{y};O)\\ 0,&\text{otherwise}\\ \end{array}\right.$
with $\sum_{i=m}^{k}q_{i,m}=1$. The distance $\Theta(\mathbf{x}_{i},C_{m};O)$ reflects the dissimilarity between $\mathbf{x}_{i}$ and $C_{m}$ computed based on the distance structures reflected by the orders $O$. Therefore, the problem to be solved can be specified as how to jointly learn $\mathbf{Q}$ and $O$ to minimize $L$. When the algorithm converges, a sub-optimal solution with $k$ compact bank client clusters is expected to be obtained. In general, “compact” means that the bank clients with more similar characteristics (i.e., with more similar value descriptions across all the attributes) are gathered into the same cluster, and different clusters are relatively distinguishable.
Frequently used symbols in this paper and the corresponding explanations are sorted out in Table 1 below. Then we discuss how to define the order distance and learn order for accurate categorical data clustering.
Table 1. Frequently used symbols. Note that we uniformly use lowercase, uppercase, bold lowercase, and bold uppercase to indicate value, set, vector, and matrix, respectively.
Symbol Explanation
$X$, $A$, $O$, and $C$ Dataset, attribute set, order set, and cluster set
$\mathbf{x}_{i}$ and $x_{i,r}$ $i$-th data object and $r$-th value of $\mathbf{x}_{i}$
$l_{r}$ Total number of possible values in $V_{r}$
$O_{r}$ Order values (integers) corresponding to $V_{r}$
$o(v_{r,g})$ Order value (integer) of possible value $v_{r,g}$
$\mathbf{Q}$ and $C_{m}$ Object-cluster affiliation matrix and $m$-th cluster
$d_{i,r,g}$ Order distance between $x_{i,r}$ and $v_{r,g}$ ($g$-th value of $\mathbf{d}_{i,r}$)
$p_{m,r,g}$ Occurrence probability of $v_{r,g}$ in $C_{m}$ ($g$-th value of $\mathbf{p}_{m,r}$)
$O_{r,m}$ Order values of $O_{r}$ obtained through $C_{m}$
$L_{r,m}$ A fraction of $L$ jointly contributed by $\mathbf{a}_{r}$ and $C_{m}$
3.2. Order Distance
We first define the sample-cluster distance $\Theta(\mathbf{x}_{i},C_{m};O)$ with considering the order $O$, to connect the order with the clustering objective. The distance $\Theta(\mathbf{x}_{i},C_{m};O)$ is defined as:
(3) $\Theta(\mathbf{x}_{i},C_{m};O)=\frac{1}{d}\sum_{r=1}^{d}\theta(x_{i,r},\textbf{p}_{m,r})$

### v3

@@
Abstract
Categorical data composed of qualitative valued attributes are ubiquitous in machine learning tasks. Due to the lack of a well-defined metric space, categorical data distributions are difficult to understand intuitively. Clustering is a popular data analysis technique suitable for understanding data distribution. However, the success of clustering often relies on reasonable distance metrics, which happen to be what categorical data naturally lack. This paper, therefore, introduces a new finding that the order relation among attribute values is the decisive factor in clustering accuracy, and is also the key to understanding categorical data clusters, because the essence of clustering is to order the clusters in terms of their admission to samples. To obtain the clustering-friendly order distance metric, we propose a new learning paradigm that allows joint learning of clusters and the orders. It alternatively partitions data into clusters based on order distance, and estimates the most likely orders according to the clusters. The algorithm achieves superior clustering accuracy with a convergence guarantee, and the learned orders facilitate the understanding of the non-intuitive categorical data from the cluster perspective. Experiments with ablation studies, significance tests, case studies, etc., have validated the efficacy of the proposed method. The source code is available at https://anonymous.4open.science/r/OCL-demo.
@@ Abstract
1 Introduction
Figure 1: Clustering accuracy of $k$-modes by treating 12 real benchmark datasets with Randomly generated Orders (RO) (color-filled regions), without orders (green bars), and with inherent semantic orders (brown bars). It can be seen that the performance upper bound of color-filled regions is obviously higher than that of the bars, indicating that even random orders can easily achieve significant accuracy improvements. Detailed statistics of the datasets and settings are provided in Table 6 and Section A.3.1 of the Appendix.
@@ 1 Introduction
In the literature11footnotetext: Data composed of nominal and ordinal attributes., clustering algorithms for categorical data can be roughly classified into: 1) $k$-modes and its subsequent variants (Huang, 1998)(Bai et al., 2011)(Jia & Cheung, 2018)(Bai & Liang, 2022) that partition data samples into clusters based on certain distance measures, and 2) cluster quality assessment-based algorithms (Hu et al., 2022)(Li et al., 2004)(Cai et al., 2024)(Zhang et al., 2025) that adjust the sample-cluster affiliations based on certain cluster quality measures. Their performance is commonly dominated by the metric space represented by the adopted measures. Thus, the focus of current research has shifted from designing clustering algorithms to proposing reasonable measures that suit the categorical data clustering (Alamuri et al., 2014).
The conventional Hamming distance measure and one-hot encoding strategy are very common in categorical data clustering, but they both simply represent the distance between two attribute values based on whether the two values are identical or not, which is far from informative. Later, on the one hand, more advanced distance measures have been proposed to discriminatively indicate the value-level differences by considering the intra-attribute probability of possible values (Sandler & Lindenbaum, 2011), inter-attribute relationship (Ahmad & Dey, 2007)(Le & Ho, 2005)(Ienco et al., 2012), or both (Jia et al., 2016)(Zhang & Cheung, 2022). On the other hand, advanced representation strategies that encode distances among data samples (Qian et al., 2015) (Zhang et al., 2023), and couplings in the attribute and value levels (Jian et al., 2018c)(Jian et al., 2017), have also been proposed. Although they have considerably improved the clustering performance, they are all task-independent and thus incompetent in adapting to different clustering tasks, i.e., $T(k,X)$ subject to different numbers of clusters $k$ on various categorical datasets $X$.
@@
It is noteworthy that order is the actual basis for clustering in most cases. For instance, when a $k$-means-type algorithm tries to partition a sample into one of the $k$ clusters, it actually obtains the overall order of the $k$ clusters in terms of admitting the sample. The overall order is formed based on the similarity orders of the clusters to the sample reflected by different attributes. Note that each numerical attribute corresponds to a one-dimensional Euclidean distance space with a clear order of its values, which makes it easy to obtain the value order. According to the above analysis, we also note that although most existing categorical data clustering approaches do not specify the order of attribute values, their clustering process still passively forms the order through similarity measurement, regardless of whether there is an obvious semantic order of the attribute values.
@@
Therefore, knowing the correct order of attribute values may considerably benefit categorical data clustering, and the demo experimental results in Figure 1 confirmed this. It can be seen that even randomly generated orders can easily achieve significant accuracy improvements. Another interesting observation is that, although attribute values of some datasets (i.e., the right eight datasets) are with obvious semantic orders, they are rarely optimal for clustering. This is because the human-assigned semantic orders are relatively subjective and may thus be independent of specific clustering tasks $T(k,X)$, preventing the semantic orders from well-serving clustering. Therefore, obtaining the optimal orders w.r.t. certain clustering tasks is the key to accurate clustering.
@@
This paper, therefore, proposes a new clustering paradigm that dynamically learns orders based on the current data partition during clustering. Since we connect the order learning to the clustering task, the optimal order for clustering can be obtained. To make full use of the learned orders, a distance learning mechanism is also introduced to dynamically measure the sample-cluster affiliation under the guidance of the orders. It turns out that the learned orders can significantly improve clustering accuracy and can also intuitively help understand the distribution and clustering effect of categorical data. Extensive experiments demonstrate the superiority of the proposed paradigm. The main contributions of this work are summarized into three points:
@@
  • A new insight that “obtaining the optimal orders is the decisive factor for accurately clustering categorical data” is proposed. Accordingly, a new clustering paradigm with order distance learning is presented.
@@
  • We propose an iterative learning mechanism to realize the optimization of the order and distance. This learning process is jointly performed with clustering, making the clustering performance more adaptive to various clustering scenarios $T(k,X)$.
@@
  • Both the learning process and learned orders are highly interpretable, which helps understand complex categorical data. Moreover, the algorithm is efficient, strictly convergent, and can be easily adapted for mixed data [^2] Dataset composed of categorical and numerical attributes. Corresponding evaluation can be found in Appendix A.10..
2 Related Work
@@
This section overviews the related works in: 1) data encoding, 2) representation learning, and 3) consideration of value order, for categorical data clustering.
@@
The data encoding techniques convert the categorical data into numerical values based on certain strategies, and then perform clustering on the numerical encoding. As the conventional one-hot encoding and Hamming distance cannot distinguish different dissimilarity degrees (dos Santos & Zárate, 2015)(Boriah et al., 2008) of data samples, information loss is usually unavoidable. Some statistic-based distance measures (Lin, 1998)(Cheung & Jia, 2013)(dos Santos & Zárate, 2015) consider the intra-attribute occurrence frequency of values to better distinguish the data samples. Some metrics (Ahmad & Dey, 2007)(Le & Ho, 2005)(Ienco et al., 2012) further define distance considering the relevance between attributes. Later, the measures (Jia et al., 2016)(Jian et al., 2018a) and encoding strategies (Jian et al., 2018c)(Jian et al., 2017)(Qian et al., 2015) that more comprehensively consider both the intra- and inter-attribute couplings reflected by the data statistics have also been presented.
@@
Recently, more advanced representation learning methods have been proposed to learn the categorical data encoding. The work proposed in (Zhu et al., 2022) projects dataset through different kernels for more informative coupling representation, and then learns the kernel parameters to obtain task-oriented representation. In contrast, (Cai et al., 2024) reveals the complex nested multi-granular cluster effect to obtain multi-granular distribution encoding representation of categorical data. Some approaches (Jian et al., 2018b)(Zhu et al., 2020) have also been designed to simultaneously encode the numerical and categorical attributes. However, the involved hyper-parameters have a great impact on their performance, and the appropriate parameter selection will be laborious in different real applications. Although a parameter-free learning approach (Jia & Cheung, 2018) has been proposed to learn sample-cluster distance and number of clusters simultaneously, the distance is updated dependent on cluster statistics and thus cannot reveal the natural distance structure of categorical data.
@@
Some new advances proposed with the consideration of value order further distinguish categorical attributes into nominal and ordinal attributes. The distance metrics (Zhang et al., 2020b)(Zhang & Cheung, 2022) define the distance between attribute values based on the data statistical information with considering the semantic order. Later, a representation learning approach was introduced by (Zhang et al., 2020a), which treats the possible values of an ordinal attribute as linearly arranged concepts, and learns the distances between adjacent concepts. Later, the works (Zhang et al., 2022b)(Zhang & Cheung, 2023) further unify the distances of nominal and ordinal attributes, and make them learnable with clustering. The most recent work (Zhang et al., 2022a) attempts to convert the nominal attributes into ordinal ones through geometric projection, and then learn the latent distance spaces during clustering. The above recent advances achieve considerable clustering performance improvements and empirically prove the importance of value order in categorical data clustering. However, they all assume the availability of semantic order of attribute values. Moreover, they directly adopt this order for clustering, and overlook that the semantic order may not be suitable under different clustering scenarios as shown in Figure 1.
3 Proposed Method
@@
This section first formulates the problem, then proposes the order distance definition, order learning mechanism, and the joint learning of the order-guided distances and clusters with time complexity and convergence analysis.
3.1 Problem Formulation
The problem of categorical data clustering with order learning is formulated as follows. Given a categorical dataset $X=\{\mathbf{x}_{1},\mathbf{x}_{2},...,\mathbf{x}_{n}\}$ with $n$ data samples. Each data sample $\mathbf{x}_{i}$ can be denoted as a $d$-dimensional row vector $\mathbf{x}_{i}=[x_{i,1},x_{i,2},..,x_{i,d}]^{\top}$ represented by $d$ attributes $A=\{\mathbf{a}_{1},\mathbf{a}_{2},...,\mathbf{a}_{d}\}$. We use $l_{r}$ to indicate the total number of possible values of an attribute, and it usually satisfies $l_{r}\ll n$ for real categorical data. Assume that there is an optimal order indicating an appropriate underlying distance structure of possible values w.r.t. each clustering task $T(k,X)$ subject to the sough number of clusters $k$ specified by the users and the characteristics of the dataset $X$. We denote the optimal order of an attribute $\mathbf{a}_{r}$ as $O^{*}_{r}=\{o^{*}(v_{r,1}),o^{*}(v_{r,2}),...,o^{*}(v_{r,l_{r}})\}$ where the superscript “*” marks the optimum and the value of $o^{*}(v_{r,g})$ is an integer reflecting the unique ranking of $v_{r,g}$ among all the $l_{r}$ possible values in $V_{r}$.
@@ 3.1 Problem Formulation
We aim to approximate the optimal orders of all the $d$ attributes $O^{*}=\{O^{*}_{1},O^{*}_{2},...,O^{*}_{d}\}$. Then the order information reflecting the distance structure can guide the partition of data objects during clustering. For a crisp partitional clustering task $T(k,X)$ to partition $X$ into $k$ non-overlapping subsets $C=\{C_{1},C_{2},...,C_{k}\}$, the objective can be formalized as the problem of minimizing:
$L(\mathbf{Q},O)=\mathop{\sum\limits_{m=1}^{k}{\mathop{\sum}\limits_{i=1}^{n}}}\cdot\Theta(\mathbf{x}_{i},C_{m};O)$ (1)
@@
$q_{i,m}=\left\{\begin{array}[]{ll}1,&\text{if}\ m=\arg\min\limits_{y}\Theta(\mathbf{x}_{i},C_{y};O)\\ 0,&\text{otherwise}\\ \end{array}\right.$ (2)
with $\sum_{i=m}^{k}q_{i,m}=1$. $\Theta(\mathbf{x}_{i},C_{m};O)$ reflects the dissimilarity (also called distance hereinafter) between $\mathbf{x}_{i}$ and $C_{m}$ computed based on the distance structures reflected by the orders $O$. Therefore, the problem can be specified as how to jointly learn $\mathbf{Q}$ and $O$ to minimize $L$. Frequently used symbols and more detailed problem settings with intuitive examples are provided in Appendix A.1. Then we present the definition of order distance $\Theta$ subject to order $O$ and discuss how to learn $O$ for categorical data clustering [^3] In the following, $\Theta$ is expressed in the form of a matrix $\mathbf{D}$ to facilitate the description of the distance definition and computation..
3.2 Order Distance
We first define the sample-cluster distance $\Theta(\mathbf{x}_{i},C_{m};O)$ with considering the order $O$, to connect the order with the clustering objective:
$\Theta(\mathbf{x}_{i},C_{m};O)=\frac{1}{d}\sum_{r=1}^{d}\theta(x_{i,r},\textbf{p}_{m,r})$ (3)